# Active Learning to Guide Labeling Efforts for Question Difficulty Estimation

### v1

@@ 1 Introduction
Question Difficulty Estimation (QDE), also known as question calibration, is a regression task that estimates a question’s difficulty directly from the question and answers’ text. It is a crucial task in personalized support tools like computerized adaptive testing Van der Linden and Glas (2000), which tailors questions to a student’s skill level. If the questions are too easy or too difficult, the student might lose motivation, negatively affecting their learning outcome Wang et al. (2014).
Traditionally, QDE has been performed with manual calibration Attali et al. (2014) and pretesting Lane et al. (2016), which are time-consuming and expensive. Recent studies aim to address these limitations by leveraging natural language processing (NLP) techniques. The NLP approaches train machine learning models to estimate question difficulty from its text. Once trained, the models can quickly calibrate unseen questions, reducing the need for pretesting and manual calibration.
Supervised techniques dominate QDE with state-of-the-art results Zhou and Tao (2020); Benedetto et al. (2021) by fine-tuning the publicly available pre-trained models BERT Devlin et al. (2018) and DistilBERT Sanh et al. (2019). However, fine-tuning often requires a large labeled dataset containing tens of thousands of calibrated questions, almost impossible to collect for individual course instructors developing QDE tools on their exam data. An isolated study Loginova et al. (2021) has delved into an unsupervised approach, relying solely on additional pre-training and evaluating pairwise difficulty. Although this approach is helpful, its performance cannot be directly compared to supervised methods and is more computationally expensive in practical implementations.
In this work, we explore active learning (AL) Settles (2009) for QDE, a data-efficient supervised approach aiming to minimize the labeling work for human annotators while matching the performance of state-of-the-art models. AL operates by iteratively training a model on an increasingly growing labeled subset by acquiring labels from an expert only for the most informative unlabeled data points. This human-in-the-loop strategy allows us to preserve the well-established supervised evaluation methods, effectively bridging the gap between the performance-driven supervised domain and the data-centric unsupervised domain. Moreover, we propose a novel acquisition function PowerVariance to add the most informative samples to the labeled set while limiting redundant information, a regression extension to the PowerBALD function Kirsch et al. (2021) popular in classification. We use DistilBERT Sanh et al. (2019) for QDE and find informative samples by applying Monte Carlo (MC) dropout Gal and Ghahramani (2016) to capture epistemic uncertainty over the unlabeled samples.
@@ 2 Related Work
Earliest NLP research on QDE from text primarily focused on multiple-choice questions (MCQs), employing bag-of-words techniques and assessing similarities among questions, correct choices, and incorrect choices Alsubait et al. (2013); Yaneva et al. (2018); Kurdi et al. (2017). However, these methods have been outperformed by more recent machine learning approaches.
Machine learning approaches to QDE fall into two main categories: (i) those involving distinct feature engineering and regression phases, and (ii) end-to-end neural networks (NNs). The former encompasses a wide range of features, including linguistic features, text embeddings, frequency-based features, and readability indexes. Several studies have also explored combinations of these feature techniques Benedetto (2023). Common machine learning regression models in this group include random forests, support vector machines, and linear regression Benedetto et al. (2023).
End-to-end NN approaches in previous research primarily rely on Transformer models Vaswani et al. (2017), which can be either supervised or unsupervised. Transformers are attention-based NNs pre-trained on a large corpus of text. This generally yields superior performance with shorter training times compared to training NNs from scratch, leveraging the pre-existing knowledge of the pre-trained model.
Supervised estimation to QDE is most prevalent in the literature Cheng et al. (2019); Qiu et al. (2019); Tong et al. (2020). Fine-tuning the publicly available pre-trained models BERT Devlin et al. (2018) and DistilBERT Sanh et al. (2019) on the task of QDE gives state-of-the-art results Zhou and Tao (2020); Benedetto et al. (2021) and has been shown to outperform other approaches using traditional NLP-derived features Benedetto (2023).
Unsupervised estimation, aiming to avoid relying on labeled data entirely, has received comparatively less attention. One study Loginova et al. (2021) estimates question difficulty by leveraging the epistemic uncertainty in question answering models as an indicator of human-perceived difficulty. This approach involves additional pre-training without fine-tuning, making it independent of labeled data. While helpful in estimating difficulty, its performance cannot be directly compared to supervised estimation as it evaluates pairwise difficulty. Moreover, it poses computational challenges in practice because numerous pairwise evaluations are required to determine an overall difficulty ranking of unseen questions.
@@ 3.1 Active Learning
AL Settles (2009) is a human-in-the-loop technique for achieving data efficiency. Instead of collecting and labeling a large dataset before training, which is time-consuming and expensive, AL iteratively acquires labels from an expert annotator only for the most informative data points from a pool of unlabeled data. After each acquisition step, the newly labeled points are added to the training set, and the model is retrained. This process is repeated until reaching a desired level of accuracy or until the labeling budget is exhausted, aiming to minimize the labeling work of human annotators. Figure 1 provides a visual overview of the AL workflow, employing pool-based sampling as described.
Figure 1: Active learning workflow with pool-based sampling. Active learning iteratively trains on a subset of labeled data and acquires labels from an expert annotator for samples in the unlabeled pool. Adapted from Settles (2009).
In AL, the informativeness of new points is assessed by an acquisition function. The acquisition function typically relies on epistemic uncertainty over unlabeled data, which can be obtained with approximate Bayesian inference techniques like MC dropout Gal and Ghahramani (2016) or with ensembling techniques Lakshminarayanan et al. (2017); Thuy and Benoit (2023, 2024). Epistemic uncertainty represents uncertainty in the model parameters and is naturally high in regions of the input space with few training observations Der Kiureghian and Ditlevsen (2009), precisely the observations we want to add to the labeled set. For classification tasks, a commonly used acquisition scoring function is Bayesian Active Learning by Disagreement (BALD) Houlsby et al. (2011), estimates the epistemic uncertainty by measuring the variability among samples of the predictive distribution. Data points maximize this acquisition function when the model assigns the highest predicted probability to a different class in each sample. For regression tasks, the epistemic uncertainty is estimated by the Variance among predictive samples Settles (2009). Similarly, data points score high on this acquisition function when the model’s output varies strongly across the samples.
@@ 3.2 Monte Carlo Dropout Uncertainty
Uncertainty in predictions can arise from two different sources: aleatoric and epistemic uncertainty Der Kiureghian and Ditlevsen (2009). Aleatoric uncertainty refers to the notion of randomness and is related to the data-measurement process. This uncertainty is irreducible even if more data is collected. Epistemic uncertainty accounts for uncertainty in the model parameters. In contrast to data uncertainty, collecting more data can reduce model uncertainty. As such, it is interesting for acquisitions functions to select the unlabeled samples with the largest epistemic uncertainty.
@@
In MC dropout Gal and Ghahramani (2016), dropout is not only applied at training time but also at test time. Multiple forward passes are performed, each time randomly dropping units and getting another thinned dropout variant of the NN. As such, it can be seen as an implicit ensemble method where each sample corresponds to an ensemble member. The various samples approximate the true posterior predictive distribution, enabling it to estimate the epistemic uncertainty in a data point.
@@ 3.3 PowerVariance Acquisition
In practical AL applications, instead of single data points, batches of data points are selected in each acquisition step to minimize the frequency of model retraining and expert involvement. A common heuristic involves selecting the top-$K$ highest-scoring points from an acquisition scheme designed for single-point selection, i.e., top-$K$ acquisition Kirsch et al. (2019) (Figure 2). However, this method overlooks interactions between points within an acquisition batch because individual points are scored independently. For example, if the most informative point is duplicated in the pool set, all instances will be acquired, which is wasteful.
@@
To address this issue, acquisition functions designed explicitly for batch acquisition with NN classifiers have been developed, such as BatchBALD Kirsch et al. (2019). These methods improve over top-$K$ acquisition by accounting for the interaction between points but are computationally expensive. To limit the computational burden, the authors of Kirsch et al. (2021) propose to stochastically acquire points following a power distribution determined by the single-acquisition scores. Intuitively, points with high acquisition scores are more likely to be sampled. For example, for BALD, the method is referred to as PowerBALD, demonstrating equal performance to state-of-the-art batch acquisition functions like BatchBALD while requiring significantly less computational resources.
The stochastic acquisition strategy Kirsch et al. (2021) assumes that as new samples are selected in a batch, future acquisition scores differ from the current scores by a perturbation. This perturbation is modeled as Gumbel-distributed noise for two reasons.
First, to select the $k$-th point in the acquisition batch of size $K$, it is important to consider how much additional information (i.e., increase in acquisition scores) the still-to-be-selected $K-k$ points will provide. As such, the stochastic strategy models the maximum future increase in acquisition scores over all possible candidate points to complete the batch. Empirically, acquisition scores are similar to a truncated exponential distribution, with different rate parameters at each AL step. The maximum over sums of such random variables is empirically shown to follow a Gumbel distribution Kirsch et al. (2021).
Second, the Gumbel distribution is also mathematically convenient. The Gumbel-Top-$K$ trick Kool et al. (2019) shows that taking the highest-scoring points from a distribution perturbed with Gumbel noise is equivalent to sampling from a softmax distribution without replacement. Building on this, perturbing the log-scores with Gumbel noise results in sampling from a power distribution. Power acquisition assumes that scores are non-negative and uninformative points should be avoided, a sensible approach for AL.
@@
Following the Gumbel-Top-$K$ trick Kool et al. (2019), taking the top-$K$ points from $s_{PowerVar}$ is equivalent to sampling without replacement from the distribution $p_{PowerVar}$
@@ 4.1 Data
RACE++ Liang et al. (2019) is a dataset of reading comprehension MCQs, built by merging the original RACE dataset Lai et al. (2017) with the RACE-C dataset Liang et al. (2019). Each question comprises a reading passage, a stem, and four possible answer options, one of them being correct. Each question has one out of three difficulty levels (0, 1, 2), which we consider as the gold standard for QDE. The difficulty levels correspond to middle school, high school, and university-level questions; the dataset is imbalanced, with a distribution of 25%, 62%, and 13% respectively. Note that the dataset labels are all available; the labels are hidden and revealed once requested by the acquisition function. The training split contains $100,568$ questions, while the validation and test splits contain $1000$ and $5642$ questions, respectively. There are no reading passages shared across the splits.
@@ 4.2 Model Architecture
We fine-tune the publicly available pre-trained model DistilBERT on the task of QDE. DistilBERT is a language model obtained by distilling BERT, i.e., compressing BERT by training a small model to reproduce its full output distribution Hinton et al. (2015). The authors of Benedetto et al. (2023) find that DistilBERT achieves comparable performance to BERT on QDE while using approximately half the parameter count. Limiting the computational expense is important in AL as the model needs to be fine-tuned over multiple iterations.
@@
Following previous research Benedetto (2023), we handle QDE on the RACE++ dataset as a discrete regression problem. The QDE model is trained as a regression model and outputs a continuous difficulty, which is then converted to the closest discrete level with simple thresholds. As evaluation metric, we compute the root mean squared error (RMSE) between the discrete predictions and discrete difficulty levels because of its consideration for the order of difficulty levels. We refer to this metric as “discrete RMSE”.
@@ 4.3 Active Learning Setup
The AL process starts with an initial labeled dataset of 500 observations, randomly selected from the training set and following the training set distribution (i.e., 25%/62%/13%). In each iteration, the model is fine-tuned for 10 training epochs and the parameters giving the best validation performance are saved. We use the AdamW optimizer Loshchilov and Hutter (2017) with learning rate 2e-5 and batch size 64.
Subsequently, the model is evaluated on a random subset of the unlabeled pool containing $5000$ observations, from which $100$ observations are selected for labeling and added to the training set. As demonstrated by Atighehchian et al. (2022), using a random subset instead of the entire pool minimally impacts predictive performance while being more computationally efficient. The configurations with MC dropout use 10 MC samples to calculate epistemic uncertainty. Before training on the new training set, the model weights are re-initialized to their original state. This involves using pre-trained weights for the base model and randomly initialized weights for the output layer with He initialization He et al. (2015), effectively mitigating the risk of getting stuck in a poor local minimum. Table 1 displays all hyperparameter settings for the model and AL setup.
@@
The experiments are implemented in PyTorch Ansel et al. (2024) using the BaaL Atighehchian et al. (2022) and HuggingFace Wolf et al. (2020) packages, executed on an NVIDIA RTX A5000 GPU. The results are averaged over five independent runs with random seeds, with a total runtime of 90 hours.
@@ 5.1 Predictive Performance
The findings reveal that fine-tuning on the initial labeled set (500 observations; 0.5%) performs exactly in between the Majority and Supervised baselines. As the labeled training set expands, we observe a decrease in discrete RMSE across all AL configurations. This finding demonstrates that DistilBERT performs well with limited labeled data, consistent with previous research Sun et al. (2019) using the non-distilled BERT model.

### v2

@@ 1 Introduction
Question Difficulty Estimation (QDE), also known as question calibration, is a regression task that estimates a question’s difficulty directly from the question and answers’ text. It is a crucial task in personalized support tools like computerized adaptive testing (Van der Linden and Glas, 2000), which tailors questions to a student’s skill level. If the questions are too easy or too difficult, the student might lose motivation, negatively affecting their learning outcome (Wang et al., 2014).
Traditionally, QDE has been performed with manual calibration (Attali et al., 2014) and pretesting (Lane et al., 2016), which are time-consuming and expensive. Recent studies aim to address these limitations by leveraging natural language processing (NLP) techniques. The NLP approaches train machine learning models to estimate question difficulty from its text. Once trained, the models can quickly calibrate unseen questions, reducing the need for pretesting and manual calibration.
Supervised techniques dominate QDE with state-of-the-art results (Zhou and Tao, 2020; Benedetto et al., 2021) by fine-tuning the publicly available pre-trained models BERT (Devlin et al., 2018) and DistilBERT (Sanh et al., 2019). However, fine-tuning often requires a large labeled dataset containing tens of thousands of calibrated questions, almost impossible to collect for individual course instructors developing QDE tools on their exam data. An isolated study (Loginova et al., 2021) has delved into an unsupervised approach, relying solely on additional pre-training and evaluating pairwise difficulty. Although this approach is helpful, its performance cannot be directly compared to supervised methods and is more computationally expensive in practical implementations.
In this work, we explore active learning (AL) (Settles, 2009) for QDE, a data-efficient supervised approach aiming to minimize the labeling work for human annotators while matching the performance of state-of-the-art models. AL operates by iteratively training a model on an increasingly growing labeled subset by acquiring labels from an expert only for the most informative unlabeled data points. This human-in-the-loop strategy allows us to preserve the well-established supervised evaluation methods, effectively bridging the gap between the performance-driven supervised domain and the data-centric unsupervised domain. Moreover, we propose a novel acquisition function PowerVariance to add the most informative samples to the labeled set while limiting redundant information, a regression extension to the PowerBALD function (Kirsch et al., 2021) popular in classification. We use DistilBERT (Sanh et al., 2019) for QDE and find informative samples by applying Monte Carlo (MC) dropout (Gal and Ghahramani, 2016) to capture epistemic uncertainty over the unlabeled samples.
@@ 2 Related Work
Earliest NLP research on QDE from text primarily focused on multiple-choice questions (MCQs), employing bag-of-words techniques and assessing similarities among questions, correct choices, and incorrect choices (Alsubait et al., 2013; Yaneva et al., 2018; Kurdi et al., 2017). However, these methods have been outperformed by more recent machine learning approaches.
Machine learning approaches to QDE fall into two main categories: (i) those involving distinct feature engineering and regression phases, and (ii) end-to-end neural networks (NNs). The former encompasses a wide range of features, including linguistic features, text embeddings, frequency-based features, and readability indexes. Several studies have also explored combinations of these feature techniques (Benedetto, 2023). Common machine learning regression models in this group include random forests, support vector machines, and linear regression (Benedetto et al., 2023).
End-to-end NN approaches in previous research primarily rely on Transformer models (Vaswani et al., 2017), which can be either supervised or unsupervised. Transformers are attention-based NNs pre-trained on a large corpus of text. This generally yields superior performance with shorter training times compared to training NNs from scratch, leveraging the pre-existing knowledge of the pre-trained model.
Supervised estimation to QDE is most prevalent in the literature (Cheng et al., 2019; Qiu et al., 2019; Tong et al., 2020). Fine-tuning the publicly available pre-trained models BERT (Devlin et al., 2018) and DistilBERT (Sanh et al., 2019) on the task of QDE gives state-of-the-art results (Zhou and Tao, 2020; Benedetto et al., 2021) and has been shown to outperform other approaches using traditional NLP-derived features (Benedetto, 2023).
Unsupervised estimation, aiming to avoid relying on labeled data entirely, has received comparatively less attention. One study (Loginova et al., 2021) estimates question difficulty by leveraging the epistemic uncertainty in question answering models as an indicator of human-perceived difficulty. This approach involves additional pre-training without fine-tuning, making it independent of labeled data. While helpful in estimating difficulty, its performance cannot be directly compared to supervised estimation as it evaluates pairwise difficulty. Moreover, it poses computational challenges in practice because numerous pairwise evaluations are required to determine an overall difficulty ranking of unseen questions.
@@ 3.1 Active Learning
AL (Settles, 2009) is a human-in-the-loop technique for achieving data efficiency. Instead of collecting and labeling a large dataset before training, which is time-consuming and expensive, AL iteratively acquires labels from an expert annotator only for the most informative data points from a pool of unlabeled data. After each acquisition step, the newly labeled points are added to the training set, and the model is retrained. This process is repeated until reaching a desired level of accuracy or until the labeling budget is exhausted, aiming to minimize the labeling work of human annotators. Figure 1 provides a visual overview of the AL workflow, employing pool-based sampling as described.
Figure 1: Active learning workflow with pool-based sampling. Active learning iteratively trains on a subset of labeled data and acquires labels from an expert annotator for samples in the unlabeled pool. Adapted from (Settles, 2009).
In AL, the informativeness of new points is assessed by an acquisition function. The acquisition function typically relies on epistemic uncertainty over unlabeled data, which can be obtained with approximate Bayesian inference techniques like MC dropout (Gal and Ghahramani, 2016) or with ensembling techniques (Lakshminarayanan et al., 2017; Thuy and Benoit, 2023, 2024). Epistemic uncertainty represents uncertainty in the model parameters and is naturally high in regions of the input space with few training observations (Der Kiureghian and Ditlevsen, 2009), precisely the observations we want to add to the labeled set. For classification tasks, a commonly used acquisition scoring function is Bayesian Active Learning by Disagreement (BALD) (Houlsby et al., 2011), estimates the epistemic uncertainty by measuring the variability among samples of the predictive distribution. Data points maximize this acquisition function when the model assigns the highest predicted probability to a different class in each sample. For regression tasks, the epistemic uncertainty is estimated by the Variance among predictive samples (Settles, 2009). Similarly, data points score high on this acquisition function when the model’s output varies strongly across the samples.
@@ 3.2 Monte Carlo Dropout Uncertainty
Uncertainty in predictions can arise from two different sources: aleatoric and epistemic uncertainty (Der Kiureghian and Ditlevsen, 2009). Aleatoric uncertainty refers to the notion of randomness and is related to the data-measurement process. This uncertainty is irreducible even if more data is collected. Epistemic uncertainty accounts for uncertainty in the model parameters. In contrast to data uncertainty, collecting more data can reduce model uncertainty. As such, it is interesting for acquisitions functions to select the unlabeled samples with the largest epistemic uncertainty.
@@
In MC dropout (Gal and Ghahramani, 2016), dropout is not only applied at training time but also at test time. Multiple forward passes are performed, each time randomly dropping units and getting another thinned dropout variant of the NN. As such, it can be seen as an implicit ensemble method where each sample corresponds to an ensemble member. The various samples approximate the true posterior predictive distribution, enabling it to estimate the epistemic uncertainty in a data point.
@@ 3.3 PowerVariance Acquisition
In practical AL applications, instead of single data points, batches of data points are selected in each acquisition step to minimize the frequency of model retraining and expert involvement. A common heuristic involves selecting the top-$K$ highest-scoring points from an acquisition scheme designed for single-point selection, i.e., top-$K$ acquisition (Kirsch et al., 2019) (Figure 2). However, this method overlooks interactions between points within an acquisition batch because individual points are scored independently. For example, if the most informative point is duplicated in the pool set, all instances will be acquired, which is wasteful.
@@
To address this issue, acquisition functions designed explicitly for batch acquisition with NN classifiers have been developed, such as BatchBALD (Kirsch et al., 2019). These methods improve over top-$K$ acquisition by accounting for the interaction between points but are computationally expensive. To limit the computational burden, the authors of Kirsch et al. (2021) propose to stochastically acquire points following a power distribution determined by the single-acquisition scores. Intuitively, points with high acquisition scores are more likely to be sampled. For example, for BALD, the method is referred to as PowerBALD, demonstrating equal performance to state-of-the-art batch acquisition functions like BatchBALD while requiring significantly less computational resources.
The stochastic acquisition strategy (Kirsch et al., 2021) assumes that as new samples are selected in a batch, future acquisition scores differ from the current scores by a perturbation. This perturbation is modeled as Gumbel-distributed noise for two reasons.
First, to select the $k$-th point in the acquisition batch of size $K$, it is important to consider how much additional information (i.e., increase in acquisition scores) the still-to-be-selected $K-k$ points will provide. As such, the stochastic strategy models the maximum future increase in acquisition scores over all possible candidate points to complete the batch. Empirically, acquisition scores are similar to a truncated exponential distribution, with different rate parameters at each AL step. The maximum over sums of such random variables is empirically shown to follow a Gumbel distribution (Kirsch et al., 2021).
Second, the Gumbel distribution is also mathematically convenient. The Gumbel-Top-$K$ trick (Kool et al., 2019) shows that taking the highest-scoring points from a distribution perturbed with Gumbel noise is equivalent to sampling from a softmax distribution without replacement. Building on this, perturbing the log-scores with Gumbel noise results in sampling from a power distribution. Power acquisition assumes that scores are non-negative and uninformative points should be avoided, a sensible approach for AL.
@@
Following the Gumbel-Top-$K$ trick (Kool et al., 2019), taking the top-$K$ points from $s_{PowerVar}$ is equivalent to sampling without replacement from the distribution $p_{PowerVar}$
@@ 4.1 Data
RACE++ (Liang et al., 2019) is a dataset of reading comprehension MCQs, built by merging the original RACE dataset (Lai et al., 2017) with the RACE-C dataset (Liang et al., 2019). Each question comprises a reading passage, a stem, and four possible answer options, one of them being correct. Each question has one out of three difficulty levels (0, 1, 2), which we consider as the gold standard for QDE. The difficulty levels correspond to middle school, high school, and university-level questions; the dataset is imbalanced, with a distribution of 25%, 62%, and 13% respectively. Note that the dataset labels are all available; the labels are hidden and revealed once requested by the acquisition function. The training split contains $100,568$ questions, while the validation and test splits contain $1000$ and $5642$ questions, respectively. There are no reading passages shared across the splits.
@@ 4.2 Model Architecture
We fine-tune the publicly available pre-trained model DistilBERT on the task of QDE. DistilBERT is a language model obtained by distilling BERT, i.e., compressing BERT by training a small model to reproduce its full output distribution (Hinton et al., 2015). The authors of Benedetto et al. (2023) find that DistilBERT achieves comparable performance to BERT on QDE while using approximately half the parameter count. Limiting the computational expense is important in AL as the model needs to be fine-tuned over multiple iterations.
@@
Following previous research (Benedetto, 2023), we handle QDE on the RACE++ dataset as a discrete regression problem. The QDE model is trained as a regression model and outputs a continuous difficulty, which is then converted to the closest discrete level with simple thresholds. As evaluation metric, we compute the root mean squared error (RMSE) between the discrete predictions and discrete difficulty levels because of its consideration for the order of difficulty levels. We refer to this metric as “discrete RMSE”.
@@ 4.3 Active Learning Setup
The AL process starts with an initial labeled dataset of 500 observations, randomly selected from the training set and following the training set distribution (i.e., 25%/62%/13%). In each iteration, the model is fine-tuned for 10 training epochs and the parameters giving the best validation performance are saved. We use the AdamW optimizer (Loshchilov and Hutter, 2017) with learning rate 2e-5 and batch size 64.
Subsequently, the model is evaluated on a random subset of the unlabeled pool containing $5000$ observations, from which $100$ observations are selected for labeling and added to the training set. As demonstrated by Atighehchian et al. (2022), using a random subset instead of the entire pool minimally impacts predictive performance while being more computationally efficient. The configurations with MC dropout use 10 MC samples to calculate epistemic uncertainty. Before training on the new training set, the model weights are re-initialized to their original state. This involves using pre-trained weights for the base model and randomly initialized weights for the output layer with He initialization (He et al., 2015), effectively mitigating the risk of getting stuck in a poor local minimum. Table 1 displays all hyperparameter settings for the model and AL setup.
@@
The experiments are implemented in PyTorch (Ansel et al., 2024) using the BaaL (Atighehchian et al., 2022) and HuggingFace (Wolf et al., 2020) packages, executed on an NVIDIA RTX A5000 GPU. The results are averaged over five independent runs with random seeds, with a total runtime of 90 hours.
@@ 5.1 Predictive Performance
The findings reveal that fine-tuning on the initial labeled set (500 observations; 0.5%) performs exactly in between the Majority and Supervised baselines. As the labeled training set expands, we observe a decrease in discrete RMSE across all AL configurations. This finding demonstrates that DistilBERT performs well with limited labeled data, consistent with previous research (Sun et al., 2019) using the non-distilled BERT model.